Find the vertical and horizontal asymptotes of the graph of f(x)=3x+1/x^2-4. plot the graph of function and asymptotes


1
Expert's answer
2021-02-25T05:05:08-0500

"\displaystyle
f(x) = \frac{3x + 1}{x^2 - 4} \\

\textsf{Vertical assymtote}:\,\, x^2 - 4 = 0\\

x = \pm 2 \\

\begin{aligned}
\lim_{x \to \infty} f(x) &= \lim_{x \to \infty}\frac{3x + 1}{x^2 - 4}\\
&= \lim_{x \to \infty}\frac{\frac{3}{x} + \frac{1}{x^2}}{1 - \frac{4}{x^2}}
&= 0
\end{aligned} \\

\textsf{Horizontal assymtote}:\,\, y = 0\\"